Consider Donald and Joe who are both 30- years of age and recently graduated with a degree in Finance. Both Donald and Joe plan to retire at age 67, and the retirement plan pays a 12 percent per annum return and is also compounded monthly. Donald plans to invest $1,000 per month beginning next month into his retirement account, while Joe shall invest $2,000 per month. Joe however does not plan to begin investing until 10 years after Donald begins to invest. How much will each of the newly grads have at retirement?
1
Expert's answer
2020-11-02T07:03:04-0500

"Future Value of Annuity = P \times \frac {(1+r)n-1} {r}"


For Donald


P= $1000

r= 12/12= 1%

n= "37\times 12=" 444


Therefore for Donald Future Value= $8,192,586


For Joe


P= $2000

r= 12/12= 1%

n= "27\times 12= 324"


Therefore for Joe Future Value= $4,825,220